how do I divide and write answer in standard form for (-2-3i)/i
Answer by ptaylor(2198)

(-2-3i)/i Multiply numerator and denominator by i and we get
(-2i-3i^2)/i^2 We know that i^2=-1, sooo
(-2i+3)/-1 =
2i-3 and standard form is a+bi, so we get:
-3+2i------------------ans
CK
-3+2i=?(-2-3i)/i multiply each term by i
-3i+2i^2=?-2-3i or
-3i-2=-2-3i
OK
Hope this helps---ptaylor